OpenStudy (anonymous):

Solve the following system of equations. 2x + 2y + z = 10 3x – y + 3z = 10 2x + 3y – 2z = 6

OpenStudy (mathstudent55):

You can use elimination. Eliminate z Multiply entire first eq by 2 and add to eq 3: 4x + 4y + 2z = 10 2x + 3y - 2z = 6 ----------------- 6x + 7y = 16 call this eq. 4 Now multiply first eq by -3 and add to eq 2 -6x - 6y - 3z = -30 3x - y + 3z = 10 ------------------- -3x - 7y = -20 call this eq. 5 Now solve eq 4 and eq 5 as a system of two equations.